Draw the alkene formed when 1-heptyne is treated with HBr in the presence of peroxide.
Answer #1

solution:
           CH3-CH2-CH2-CH2-CH2-C≡CH +HBr -----> CH3-CH2-CH2-CH2-CH2-CBr=CH2

CH3-CH2-CH2-CH2-CH2-CBr=CH2 + HBr ------> CH3-CH2-CH2-CH2-CH2-CBr2-CH3

In presence peroxide,

CH3-CH2-CH2-CH2-CH2-C≡CH + HBr -----> CH3-CH2-CH2-CH2-CH2-CH=CHBr

CH3-CH2-CH2-CH2-CH2-CH=CHBr + HBr ----> Two isomeric compounds with equal possibilities of Br atom to be attached to C1 and C2.
Answer #2
CH3-CH2-CH2-CH2-CH2-C=CH +HBr -----> CH3-CH2-CH2-CH2-CH2-CBr=CH2

CH3-CH2-CH2-CH2-CH2-CBr=CH2 + HBr ------> CH3-CH2-CH2-CH2-CH2-CBr2-CH3

In presence peroxide,

CH3-CH2-CH2-CH2-CH2-C=CH + HBr -----> CH3-CH2-CH2-CH2-CH2-CH=CHBr

CH3-CH2-CH2-CH2-CH2-CH=CHBr + HBr ----> Two isomeric compounds with equal possibilities of Br atom to be attached to C1 and C2.
Answer #3
CH3-CH2-CH2-CH2-CH2-C=CH +HBr -----> CH3-CH2-CH2-CH2-CH2-CBr=CH2

CH3-CH2-CH2-CH2-CH2-CBr=CH2 + HBr ------> CH3-CH2-CH2-CH2-CH2-CBr2-CH3

In presence peroxide,

CH3-CH2-CH2-CH2-CH2-C=CH + HBr -----> CH3-CH2-CH2-CH2-CH2-CH=CHBr

CH3-CH2-CH2-CH2-CH2-CH=CHBr + HBr ----> Two isomeric compounds with equal possibilities of Br atom to be attached to C1 and C2.